OpenStudy (bookworm14):

Bruno will spin the arrow on the spinner 3 times. What is the probability that the arrow will stop on A, then B, then C? The spinner is labeled A through E

OpenStudy (ybarrap):

 The probabilities for each area are each equally likely: $$ P(A)=P(B)=P(C)=P(D)=P(E)=\cfrac{1}{5} $$ Note that \(\Large \Sigma_{i\in \text{(A-E)}}P(i)=1\) Each stop is independent of the other. So the probability of stopping on A then B then C is: $$ P(A)\times P(B)\times P(C)=\cfrac{1}{5}\times\cfrac{1}{5}\times\cfrac{1}{5}\\ =\cfrac{1}{5^3}\\ $$

OpenStudy (bookworm14):

thank you very much!